OpenStudy (anonymous):

We can see if a number is divisible by 9 by adding up the digits. Can we prove that this works?

OpenStudy (anonymous):

if the sum of the digits is a multiple of 9, then the number is divisible by 9. take 81 for example, 81/9 = 9, 8+1 =9. another example would be 99, 99/9 = 11, 9 + 9 = 18 = 9*2

OpenStudy (anonymous):

is that a proof though?

OpenStudy (anonymous):

somewhat, matters on sum.

OpenStudy (anonymous):

I'm not good with formal proofs...This would be my work for the proof: let x represent the sum of individual digits of a number if x is divisible by 9, then the number is divisible by 9. if not, then the number is not divisible by 9.